Differentiate between strong convergence of a random variable and weak convergence in probability


1
Expert's answer
2021-10-12T10:30:31-0400

Solution:

For weak convergence law:

"\lim _{n \rightarrow \infty} P\left(\left|Y_{n}-Y\right|<\epsilon\right)=1"

For strong convergence law:

"P\left(\lim _{n \rightarrow \infty} \| Y_{n}-Y||<\epsilon\right)=1"